Ship reports banner with links Home IODP-USIO daily/weekly reports

< October | Next expedition >

IODP Expedition 336

Mid-Atlantic Ridge Microbiology

Daily Science Report for 16 November 2011

LOCATION: Ponta Delgada, Azores

SCIENCE UPDATE: Arrived Ponta Delgada, Azores at 0836. Successfully concluded IODP Mid-Atlantic Microbiology Expedition 336. We installed two CORK observatories in oceanic crust, prepared a third borehole for installation of an observatory to be installed in April 2012, cored 362 m of basement and recovered 82 m at two sites (which were sampled for microbiology), collected downhole logs of oceanic crust at three sites (including a novel microbiology logging tool), and cored the overlying sediment and basement contact in four holes, which have been intensively sampled for microbiology and porewater geochemistry.


Daily Science Report for 15 November 2011

LOCATION: In transit to Ponta Delgada, Azores

SCIENCE UPDATE: All science reports are completed. Scientists continued their postcruise research discussions and submitted their draft titles for postcruise research/publications. Completed lab cleaning and final preparation of shipments. Estimated time of arrival at pilot station is 0730 hr tomorrow (16 November).


Daily Science Report for 14 November 2011

LOCATION: In transit to Ponta Delgada, Azores

SCIENCE UPDATE: Scientists are preparing final reports, final samples for shipment, and discussing postcruise research collaborations. Technical staff is preparing for port call including preparing all shipments, lab reports, and closing down/cleaning laboratories. We are now confirmed to arrive at Ponta Delgada on 16 November.


Daily Science Report for 13 November 2011

LOCATION: In transit to Ponta Delgada, Azores

SCIENCE UPDATE: Scientists and Technical Staff are completing a few final analyses as well as preparing final reports and samples for shipment. Our transit speed has been better than expected – current ETA in Ponta Delgada is 16 November pending availability of a berth 1 day earlier than originally planned.


Daily Science Report for 12 November 2011

LOCATION: In transit to Ponta Delgada, Azores

SCIENCE UPDATE: Today we finished processing the whole-round cores through the track systems and all cores have been split, described, and sampled for physical properties; a few pore water analyses are ongoing. Scientists and technical staff are preparing samples for shipment.


Daily Science Report for 11 November 2011

LOCATION: In transit to Ponta Delgada, Azores

SCIENCE UPDATE: After completing drilling operations at 0045, we started the 1385 nmi transit to Ponta Delgada, Azores. Our estimated time of arrival is 0730, 17 Nov. Processing, sampling, and sample analyses of the sediment cores are ongoing. End of expedition activities are in full swing.


Daily Science Report for 10 November 2011

LOCATION: Hole U1384A (NP-1; 22.8118°N, 46.0891°W, water depth 4476 mbrf)

SCIENCE UPDATE: Today we finished coring in Hole U1384A. Cores U1384A-3H to 11H penetrated from 14.1 to 94.7 mbsf and recovered 78.87 m (98%) of nannofossil foraminifer ooze. We then penetrated 1.5 m into the underlying basalt (Core 11X) and recovered 0.58 m. We continue to intensively sample the cores for microbiologic and geochemical studies – today's totals are 975 from Hole U1384A. After coring was completed at 1255, we started retrieving the drill string and preparing for the transit to the Azores.


Daily Science Report for 9 November 2011

LOCATION: Hole U1382B (Hole 395A; 22.7559°N, 46.0812°W, water depth 4494 mbrf)
and Hole U1384A (NP-1; 22.8118°N, 46.0891°W, water depth 4476 mbrf)

SCIENCE UPDATE: We finished coring in Hole U1382B with Cores 6H to 10H (43.5 to 90 mbsf) recovering 48.92 m (105%) of nannofossil foraminifer ooze. We then penetrated 8 m into the underlying basalt (Cores 11X and 12X) and recovered 0.58 m. In the afternoon, we moved to Hole U1384A (NP-1) and started APC coring. Cores U1384A-1H and 2H were taken from 0 to 14.1 mbsf and recovered 14.64 m (103%) of nannofossil foraminifer ooze. We continue to intensively sample for microbiologic and geochemical studies – today’s totals are 615 from Hole U1382B and 165 from Hole U1384A. Tomorrow, we anticipate finishing Hole U1384A. Any time remaining before departure will be spent APC coring the lowermost sediments above basement at Sites NP-12 and -11.


Daily Science Report for 8 November 2011

LOCATION: Hole U1383E (NP-2; 22.8021°N, 46.0526°W, water depth 4425 mbrf)
and Hole U1382B (Hole 395A; 22.7559°N, 46.0812°W, water depth 4494 mbrf)

SCIENCE UPDATE: Today we cored from 0 to 44.2 m in Hole U1383E – our second APC sediment hole at Site U1383. Cores U1383E-1H to 6H recovered 49.98 m of nannofossil foraminifer ooze. Core 7X extended 1 m into basement and recovered 0.31 m of basalt rubble. In the middle of the day, we moved back to Site U1382 and started APC coring in Hole U1382B. Cores U1382B-1H to 5H extend from 0 to 43.5 mbsf and recovered 34.78 m of nannofossible foraminifer ooze. We have approximately four more APC cores until we hit basement. Many samples were taken of the sediment overlying basement for intensive microbiologic and geochemical studies – 675 from Hole U1383E and over 400 from Hole U1382B so far. Tomorrow, we anticipate finishing Hole U1382B and starting Hole U1384A (NP-1).


Daily Science Report for 7 November 2011

LOCATION: Hole U1383D (NP-2; 22.8019°N, 46.0527°W, water depth 4425 mbrf)

SCIENCE UPDATE: Today we cored from 0 to 44.3 m in Hole U1383D (NP-2). Cores 1H to 6H recovered 47.71 m of nannofossil foraminifer ooze. Core 7X extended 1 m into basement and recovered 0.76 m of basalt. Just around the midnight hour, we were offsetting to core one final APC/XCB hole at this site. Nearly 700 samples have already been taken of the sediment overlying basement for intensive microbiologic and geochemical studies.


Daily Science Report for 6 November 2011

LOCATION: Hole U1383C (NP-2; 22.8021°N, 46.0528°W, water depth 4425 mbrf)

SCIENCE UPDATE: Today we completed the installation of the 247.6 m long, multi-level CORK observatory in Hole U1383C. After deployng the ROV platform over the CORK onto the reentry cone, we released the running tool from the CORK head at 2000 hr. This completed the North Pond Expedition (336) observatory operations. During the four days remaining before we depart for the Azores, we will core the sediment overlying basement for intensive microbiologic and geochemical studies.


Daily Science Report for 5 November 2011

LOCATION: Hole U1383C (NP-2; 22.8021°N, 46.0528°W, water depth 4425 mbrf)

SCIENCE UPDATE: After recovering the logging/packer bottom-hole assembly at 0300, the rest of the day was spent assembling the 247.6 m long CORK observatory. This includes three packer assemblies to isolate three separate depth intervals in oceanic crust. Five external umbilicals allow access to these three zones for microbiological, geochemical, pressure sampling. The CORK tubing (coated steel and fiberglass) extends to 247.6 mbsf and includes perforated and slotted intervals (67.4 to 129.4, 154.8 to 181.1 mbsf, and 203.7 to 246.6 mbsf) that provide access to the three isolated intervals. An internal OsmoSampler string extends the full length of the CORK, includes seals isolating the three zones and microbiological, geochemical, and temperature experiments.


Daily Science Report for 4 November 2011

LOCATION: Hole U1383C (NP-2; 22.8021°N, 46.0528°W, water depth 4425 mbrf)

SCIENCE UPDATE: Today was spent conducting hydrologic (packer) experiments of the oceanic crust. Our plan was to conduct multiple flow tests at three different levels in the borehole. The first level was testing the entire hole by setting the packer in the lowermost portion of the casing. Due to heave and the smoothness of the interior of the casing, we could not get the packer to set. We successfully completed a full test in the open hole with the packer set at ~143 mbsf (~105 m into basement). The packer would not inflate at the deepest packer seat and it was as recovered damaged, perhaps during the first test in open hole. After the packer test ended, we lowered the logging bit to the bottom of the hole to make sure it's ready for receiving the CORK. The rest of the day was spent (1) dropping an ROV platform over the previous hole's reentry cone (Hole U1383B), (2) retrieving the logging drill string, and (3) preparing the moon pool for assembling the CORK.


Daily Science Report for 3 November 2011

LOCATION: Hole U1383C (NP-2; 22.8021°N, 46.0528°W, water depth 4425 mbrf)

SCIENCE UPDATE: Today we successfully completed wireline logging with two different tool strings. The first tool string included the new in situ tool for detecting microbial life in ocean floor boreholes - the Deep Exploration Biosphere Investigative tool (DEBI-t). The Formation MicroScanner and Sonic tool string was the second tool string. Both strings were able to collect data over the entire borehole. These are providing critical information for designing the subsequent hydrologic (packer) experiment and CORK observatory installation. At the end of the day, we were preparing to initiate the hydrologic experiment.


Daily Science Report for 2 November 2011

LOCATION: Hole U1383C (NP-2; 22.8021°N, 46.0528°W, water depth 4425 mbrf)

SCIENCE UPDATE: We finished preparing the hole for logging, hydrologic (packer) experiments, and CORK installation. Multiple tight spots and 4 m of fill on the bottom were observed on the first of three wiper trips. After two more wiper trips that were able to get back to bottom without observing any tight spots or fill, we pulled out of the hole. Once the bit was back on board, we assembled and painted the perforated and coated drill collars for the lowermost part of the CORK to be deployed a few days from now. The rest of the day was spent assembling the logging and packer bottom hole assembly and lowering it to the seafloor.


Daily Science Report for 1 November 2011

LOCATION: Hole U1383C (NP-2; 22.8021°N, 46.0528°W, water depth 4425 mbrf)

SCIENCE UPDATE: Today we finished RCB coring in Hole U1383C. The last three cores (30R-32R) were taken from 303.7 to 331.5 mbsf (265.4 to 293.2 m below the sediment/basement interface) and recovered 7.82 m of basalt (28%). After lunch we started cleaning the hole in preparation for logging, hydrologic (packer) experiments, and CORK installation. We expect to start logging late tomorrow.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Thursday, 17-Nov-2011 07:38:09 CST.